Q2) What is the pH at which a buffer composed of CO32- and HCO3- would be most effective at resisting pH change? Round your answer to 2 decimal places. Remember you can find KA and/or KB values in your textbook
Answer #1

Sol.

As Ka of HCO3- = 4.8 × 10-11  

So , pKa = - log(Ka) = - log( 4.8 × 10-11 ) = 10.32  

As Henderson - Hasselbalch equation ,  

pH = pKa + log ( [CO32-] / [HCO3-] )  

pH would be most effective , when , the concentration of conjugate base , CO32- and concentration of acid , HCO3- are equal  

That is , [HCO3-] = [CO32-]

So , pH = pKa + log ( 1 ) = pKa   

pH =   10.32